OpenStudy (anonymous):

\[v^2-2v-35/v^2-10v+25 * v-5/v-7\] Multiply and simplify

OpenStudy (anonymous):

is it??? v² - 2v - 35 v - 5 ------------ * ------ v² - 10v + 25 v - 7

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

v² - 2v - 35 v - 5 v² -7v + 5v -35 v - 5 ------------ * ------ = --------------- * ------ v² - 10v + 25 v - 7 v² - 5v -5v + 25 v - 7 v² -7v + 5v -35 v - 5 v(v-7) +5(v-7) v - 5 = --------------- * ------ = --------------- * ------ v² - 5v -5v + 25 v - 7 v(v-5) -5(v-5) v - 7 (v+5)(v-7) v-5 v + 5 = ----------- * ----- = ------- (v-5)(v-5) v-7 v - 5

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

where is my medal???? ☺

OpenStudy (anonymous):

sorry I just got interrupted. I apologize. I promise I ALWAYS give medals.

OpenStudy (anonymous):

I really am sorry I didn't give your medal right away!

OpenStudy (anonymous):

no problem...............as they say, "better late than never!!!!" ☺☻☺

OpenStudy (anonymous):

thanks. Sorry I work full time at a company and at home, go to school and have 2 little rugrats running around and it is very easy for me to get interrupted. Thank you for understanding and I will never let it happen again! :)

OpenStudy (anonymous):

even if u do not give a medal, u r always welcome to my help..... :D

OpenStudy (anonymous):

awww thanks! I like how you show the steps. Even if I still do not understand I can see how you got your answer and I can attempt to understand and I appreciate that!

OpenStudy (anonymous):

yes, my intention is to "teach" and not just give answer to get a medal.....

OpenStudy (anonymous):

:)